Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day04


Crew 101 Log Book for March 15, 2011

Photos of the day





Commander's Check-In Report

Jon Kosh Reporting

Crew Health: Mostly good.One crew is a bit under the weather, feeling colder than the rest of the crew as we sit in t-shirts. Audible shivering breaths taken.

EVAs and Times:

4: 1300 - 1315

5: 1347 - 1419

Number of Reports: 5 (Commander, Engineering, Journalist, Chef, 2 EVA)

Field Results: Geological samples taken from 3 different sites. Coordinates taken for two of the three (battery death prevented the third)

Plans for Tomorrow: Start heart rate monitor experiment, perform analysis of geological sample returns, and continue geological EVA sampling.


Commander's Report

Jon Kosh Reporting

This morning the crew awoke seemingly more comfortable with their new home, ready to begin a new sol on the Red Planet. Jon and Mitesh had radio communications fully prepared for the planned expedition and experiment specific equipment preparations went smoothly. Tommy and Christina suited up, depressurized and moved to the vehicular landing area.

After attempting to start all 4 ATV's, and following several minutes of technical diagnosis, the EVA team entered the airlock much sooner than expected. Temporarily extending SIM boundaries to include the ATV landing area, engineering crewmembers were sent to investigate. After

another round of evaluation, the crew realized more specific knowledge and expertise was required to surpass this off-nominal technical challenge and promptly notified mission support.

The vehicular sampling EVA quickly evolved into a walking distance EVA that resulted in sample returns from three different geological sites. Radio contact was maintained for the expedition and a time based event log was kept detailing GPS coordinates and accuracies as well as site

descriptions. This data log will grow to include sample analysis and visual evidence of sample sites.

The crew spent the next part of the afternoon relaxing with more episodes of the Twilight Zone and Firefly while enjoying a nice warm bowl of clam chowder. The crew commander passed on the report to the executive officer. Today is a cooking day, and just before heading in for some rest, the commander kneaded some fresh bread and set it out to soak in sunlight and rise. The crew is currently enjoying the fruits of this labor along with what smells like a fantastic spaghetti dinner. While still typing the report at the dinner table, the XO is glad to report that the crew commander seems to still have a healthy appetite. The crew's vehicular EVA's are on hold until the crew is able to receive technical consultation from mission support-which we have received word may happen tonight or tomorrow morning.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level:

o V everstart

o V armor plate

o V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): 1 generator run


 Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75

Notes/Comments:


Oil:

• Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

 engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:

 

 Water Status:

• Outside Potable Water Tank Level (inches from bottom): 11

• Trailer Potable Water Tank Level (inches from bottom): 19

• Hab Potable Water Tank Level (inches from bottom): 9

• Potable Water Meter Reading: 435396

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: a little more duckweed, level unchanged

• Condition of plants in Tank 3: water level has dropped about an inch, some plants in the back of the tank are dead and were removed

 from the tank

• Greenhab notes/ comments:

 Internet still bad. Cannot attach pictures.

 

Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:

 

Transportation:

Hab Car (NOT RENTAL)

• Used (yes/no): No

• Oil Checked (yes/no): No

• Tire Status: Not checked

• Odometer at end of day: 123688.4

• Notes/Comments on Hab Car:


 ATV 1 (Opportunity):

• Used: yes

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments: started, but not really driven since we couldn’t

 start other ATVs

 

ATV 2 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

 

ATV 3 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

 

ATV 4 (Spirit):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

 

ATV Comments: completed training on Saturday for ATVs, but don’t know how to start any but Opportunity

 

Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 68

• Actual temperature upstairs (Farenheit): 76

• Thermostat setting downstairs (Farenheit): 55

• Actual temperature downstairs (Farenheit): 65

Notes/Comments:

 

Computers/Networking Infrastructure: working better, but still goes out occasionally

 

Notes/comments on internet/computers: we have attempted to cut down even more on internet usage, and it appears to be working. Still don’t know how so much bandwidth could have been used before, but oh well.


General Engineering

Notes/Comments:


EVA 4 Report

Christina Graves Reporting

Crew Members: Commander Tom Caillouet and Christina Graves

EVA Time: 1300-1315

EVA Procedure: Commander Caillouet and I went out at 1300 to board the ATV's in order to collect samples for the geography experiment.

EVA Results: Due to mechanical failure (or human error) in the operation of the ATV's, both of us were not able to utilize the ATV's for the EVA, therefore the vehicular EVA was canceled.

Lessons Learned: Ensure all crew members know how to trouble shoot the ATV's prior to mission.



EVA 5 Report

Christina Graves Reporting

Crew Members: Commander Tom Caillouet and Christina Graves

EVA time: 1347-1419

EVA Procedure: Commander Caillouet and I set out on foot to 3 predetermined sites for soil sample collection.  At each site, a 6' x 6' area was marked, photographed, and soil samples were collected.

Each site's location was also recorded by GPS. The soil samples were set aside for analysis on 3/16/11.

EVA Results: Soil sample sites marked and photographed, and soil samples all collected successfully.

Lessons learned: none to report.


Journalist Report

Christine Redmond Reporting

Last night was close to a tragedy for any Tech student. The Hab lost all internet connection!  We ruffed through it though. I repeat crew 101 will do anything for the advancement of human space exploration! The outage lasted approximately 1.5 hours. The outage occurred exactly at the time of our mission support window for submitting our daily reports. Crew 101 managed. We notified mission support immediately of the outage and they were able to get it back up and working fairly quickly so we could submit our reports before calling it a night.

This morning was Christmas morning for Christina! We woke up to find the final shipment of our lab equipment sitting in the air lock. Christina was extremely excited to finally get all of her biology equipment. Christina and Tommy headed out for our Geology expedition mid morning. We ran into another road block. We could not get the ATVs to start for this EVA. So Christina and Tommy headed out on foot. This EVA involved observing the terrain of three separate 6 foot square areas and taking samples. The ATVs failing to start was another unanticipated problem to add to our list of many after only two days of sim. We have all learned these past few days that on Mars things don’t always go as planned and you need to be able to work with the limited resources you have to ensure a successful mission.

Today is a cook day! Tommy baked bread this afternoon. Christina and Mitesh are busy making pasta right now. We are all looking forward to a pasta dinner and a busy day tomorrow!